OpenStudy (anonymous):

David will rent a car for the weekend. He can choose one of two plans. The first plan has an initial fee of $51.96 and costs an additional $0.15 per mile driven. The second plan has an initial fee of $43.96 and costs an additional $0.20 per mile driven. How many miles would David need to drive for the two plans to cost the same?

OpenStudy (anonymous):

@ksanka

OpenStudy (triciaal):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

To find the relationship between miles and money, determine which value is constant at first, then determine which value is depended on miles the 1st plan $51.96 + $0.15*n, where n is the number of miles the 2nd plan $43.96 + $0.20*n

OpenStudy (anonymous):

now u need to find at what n the 1st and 2nd plan cost the same 51.96 + 0.15n = 43.96 + 0.20n isolate the variable n and solve

OpenStudy (anonymous):

So combine on both sides?

OpenStudy (anonymous):

subtract 43.96 from both sides and subtract 0.15n from both sides

OpenStudy (anonymous):

then combine like terms

OpenStudy (anonymous):

51.96-43.96+0.15n-0.15n = 43.96 - 43.96 + 0.20n - 0.15n

OpenStudy (anonymous):

i got 160

OpenStudy (anonymous):

8=0.05n yes n=160